### v1

# Advantage Shaping as Surrogate Reward Maximization:

Abstract
This note reconciles two seemingly distinct approaches to policy gradient optimization for the Pass@K objective in reinforcement learning with verifiable rewards: (1) direct REINFORCE-style methods, and (2) advantage-shaping techniques that directly modify GRPO. We show that these are two sides of the same coin. By reverse-engineering existing advantage-shaping algorithms, we reveal that they implicitly optimize surrogate rewards. We specifically interpret practical “hard-example up-weighting” modifications to GRPO as reward-level regularization. Conversely, starting from surrogate reward objectives, we provide a simple recipe for deriving both existing and new advantage-shaping methods. This perspective provides a lens for RLVR policy gradient optimization beyond our original motivation of Pass@K.
1 Introduction
When evaluating large language models on math and coding tasks, standard practice is to generate $K$ independent solutions and check whether any succeeds; the so-called Pass@K metric [Kul+19, Che+21]. Yet most policy gradient methods, from REINFORCE [Wil92] to more recent variants like RLOO [KHW19, Ahm+24] and GRPO [Sha+24], optimize for single-attempt reward creating a mismatch. Should we instead train what we test?
Recent work has approached this question from two seemingly distinct angles. [Mah+25, WK25] derive policy gradients that directly maximize Pass@K reward, yielding REINFORCE-style updates with reweighted advantages. Meanwhile, [Che+25] propose modifying GRPO through advantage shaping, that is, directly adjusting advantage scores to account for the Pass@K objective. Both improve Pass@K performance, but their relationship remains unclear. Are these fundamentally different techniques, or two views of the same underlying principle?
This note unifies advantage shaping and direct optimization, showing they are two sides of the same coin. As a byproduct, we establish reward regularization as a design mechanism for policy gradient optimization with verifiable rewards, providing a recipe for algorithm interpretation via reverse-engineering and development via forward-engineering.
1.1 Problem Setup: Reinforcement Learning with Verifiable Rewards
We study reinforcement learning with verifiable rewards (RLVR). We are given a distribution $\mathcal{P}$ over problem-answer pairs $(x,a)$, where $x$ might be a math problem and $a$ its final numeric answer. The goal is to train a large language model (LLM) to maximize the reward of finding correct answers to problems from $\mathcal{P}$.
We abstract the LLM as a conditional distribution $\pi_{\theta}(\text{response}|\text{prompt})$ parameterized by $\theta$. Given a prompt $x$, the model generates a response $y$ (a sequence of tokens) by autoregressive sampling. We seek parameters $\theta$ that maximize the expected reward:
$\max_{\theta}\;\;\,\mathbb{E}_{{(x,a)\sim\mathcal{P}}}\Big[\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}\,r(y,a)\Big]\,$,
where the reward function $r(y,a)$ assesses whether the model response $y$ agrees with the reference answer $a$. RLVR assumes this can be verified externally. When $y$ agrees with $a$, we call the response correct; otherwise it is wrong. Note the nested expectation: first over problem-answer pairs $(x,a)\sim\mathcal{P}$, then over model responses $y\sim\pi_{\theta}(\cdot|x)$ for each problem.
1.1.1 The 0/1 Reward Baseline
The standard reward function is the binary 0/1 reward:
$r_{0/1}(y,a):=\begin{cases}1&\text{if $y$ agrees with $a$}\\ 0&\text{otherwise}\end{cases}\,.$
Model performance is measured by the expected 0/1 reward over all examples $(x,a)$ in the distribution: $\mathbb{E}_{(x,a)\sim\mathcal{P}}[\rho_{\theta}(x,a)]$, where the per-example expected 0/1 reward (equal to the probability that model $\theta$ generates a correct response for problem $x$) is:
$\rho_{\theta}(x,a):=\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}\,r_{0/1}(y,a)\,.$
A popular algorithm for maximizing expected 0/1 reward is REINFORCE Leave-One-Out (RLOO) [KHW19, Ahm+24], which extends the classical REINFORCE algorithm [Wil92] using multiple generated responses. At each iteration $t$, and for each pair $(x,a)$ from a training set $\mathcal{D}$, RLOO updates the model parameters as
$\theta_{t+1}=\theta_{t}+\eta\cdot\widehat{G}_{\text{RLOO}}(\theta_{t};(x,a))\,$,
where, $\eta$ is the learning rate, and the gradient estimate is a weighted average over $N$ generated responses $y_{i}\stackrel{{\scriptstyle\text{IID}}}{{\sim}}\pi_{\theta}(\cdot|x)$:
$\widehat{G}_{\text{RLOO}}(\theta;(x,a))=\frac{1}{N}\sum_{i\in[N]}A^{\text{RLOO}}_{i}\cdot\nabla_{\theta}\log\pi_{\theta}(y_{i}|x)\,.$ (RLOO)
The weights $A^{\text{RLOO}}_{i}$ are called advantage scores. The construction of RLOO’s advantages scores ensures $\widehat{G}_{\text{RLOO}}$ is an unbiased, low-variance estimate of $\nabla_{\theta}\rho_{\theta}(x,a)$, the gradient of the population (expected) 0/1 reward. Intuitively, for large $N$:
$\widehat{G}_{\text{RLOO}}(\theta;(x,a))\approx\nabla_{\theta}\rho_{\theta}(x,a)\,.$ (1)
Despite its simplicity, RLOO is a strong baseline [Ahm+24]. Its structure also underlies recent variants like GRPO [Sha+24]. As noted by [Den+25], in full online mode (when clipping can be ignored), GRPO uses the same form as (RLOO) with normalized advantages $A_{i}^{\text{GRPO}}=A_{i}^{\text{RLOO}}/\sqrt{\hat{\rho}(1-\hat{\rho})}$, where $\hat{\rho}$ is the empirical 0/1 reward for example $(x,a)$.
1.1.2 The Pass@K Reward
The 0/1 reward, however, misaligns with practical LLM usage. Given a problem $x$, users typically prompt the model multiple times (say, $K$ times) and examine the responses $y_{1},\ldots,y_{K}$ to find one that produces the desired answer. Accordingly, LLMs are often evaluated using the Pass@K metric [Che+21], which measures whether at least one solution is correct among $K$ generated responses:
$r_{\text{\tiny Pass@K}}(\{y_{i}\}_{i\in[K]},a)$ $=\begin{cases}1&\text{if at least one $y_{i}$ agrees with $a$}\\ 0&\text{otherwise}\end{cases}\,.$
Since test-time performance is measured via Pass@K, a natural question arises: why not train the model to directly maximize Pass@K reward rather than 0/1 reward? That is, the learning goals becomes maximizing $\mathbb{E}_{(x,a)\sim\mathcal{P}}[\rho_{K,\theta}(x,a)]\,$, where we define the per-example expected Pass@K reward as
$\rho_{K,\theta}(x,a):=\mathbb{E}_{\{y_{i}\}_{i\in[K]}\sim\pi_{\theta}(\cdot|x)}\,r_{\text{\tiny Pass@K}}(\{y_{i}\}_{i\in[K]},a)\,.$
This equals the probability that model $\theta$ generates at least one correct response in $K$ samples.
1.2 Contributions
C-1. Direct Pass@K optimization. Building on [Mah+25], we show that the exact Pass@K policy gradient of $\rho_{K,\theta}=\rho_{K,\theta}(x,a)$ equals the 0/1 gradient reweighted by per-example Fail@(K-1) probabilities. This yields unbiased Monte Carlo estimators REINFORCEK/RLOOK using leave-one-out weights, and biased variants GRPOK. All methods naturally upweight rare successes and can be cast as asymmetric reweightings of their 0/1-reward analogues.
C-2. Advantage shaping == surrogate reward maximization. We show that the GRPO-based approach of [Che+25] can be reverse-engineered to reveal that it asymptotically (in the number of generated responses) maximizes a surrogate reward $\frac{2}{K}\arcsin(\sqrt{\rho_{K,\theta}})$. We also provide forward engineering that reproduces advantage-shaping heuristics in three steps:
  1. (i)
    Differentiate the surrogate reward using the log-derivative trick;
  2. (ii)
    Substitute all population quantities with their empirical analogues;
  3. (iii)
    Replace the population reward gradient with its RLOO proxy (Eq. (1)).
As a corollary, when K=1K{=}1, vanilla GRPO is (up to min-clip operations) effectively RLOO applied to the surrogate reward $2\arcsin(\sqrt{\rho_{\theta}})$, a variance-stabilizing transformation of the 0/1 reward.
C-3. Reward-level regularization. We interpret commonly used heuristics that downweight “easy” examples (that is, those already solved with high probability) and upweight “hard” ones [Che+25, Tea+25]. We show that multiplying GRPO gradient contributions by factors such as $1-\widehat{\rho}$ is effectively equivalent to optimizing a regularized surrogate reward $\arcsin(\sqrt{\rho_{\theta}})+\sqrt{\rho_{\theta}(1-\rho_{\theta})}$. This can be interpreted as a special instance of regularized surrogate reward maximization:
$\max_{\theta}\;\;\widehat{\mathbb{E}}_{(x,a)\sim\mathcal{D}}\left[\,F\left(\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}\,r(y,a)\right)+\lambda\cdot\Omega\left(\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}\,r(y,a)\right)\,\right]\,$, (2)
where the stochastic training objective over examples $(x,a)\sim\mathcal{D}$ consists of a data-fitting term $F$, that a strong baseline like GRPO would directly optimize, and, an additive regularizer $\Omega$ (balanced via $\lambda\geq 0$). Unlike typical policy-level or parameter-level regularization, this operates simply at the reward level. We demonstrate the regularization perspective by deriving entropy-regularized variants using the three-step recipe outlined in Contribution C-2, yielding new tunable advantage-shaping algorithms.
Notation.
For a given example $(x,a)$, let $y_{1},\ldots,y_{N}$ denote $N$ responses generated IID from the model $\pi_{\theta}(\cdot|x)$. Let $r_{i}:=r_{0/1}(y_{i},a)$ denote the 0/1 reward of response $i$, and let $\widehat{\rho}:=\frac{1}{N}\sum_{i\in[N]}r_{i}$ denote the empirical 0/1 reward. Let $N^{+}:=\sum_{i\in[N]}r_{i}$ and $N^{-}:=N-N^{+}$ denote the number of correct and wrong responses, respectively. We use the shorthand $\nabla_{i}:=\nabla_{\theta}\log\pi_{\theta}(y_{i}|x)$ for the log-probability gradient of individual response $y_{i}$, and for the average log-probability gradient over correct/wrong responses:
$\widehat{\nabla}_{+}:=\frac{1}{N^{+}}\sum_{i:r_{i}=1}\nabla_{\theta}\log\pi_{\theta}(y_{i}|x)\,\quad\text{and}\quad\widehat{\nabla}_{-}:=\frac{1}{N^{-}}\sum_{i:r_{i}=0}\nabla_{\theta}\log\pi_{\theta}(y_{i}|x)\,.$ (3)
2 Warm-up: Optimizing the 0/1 Reward
With access to a finite training set $\mathcal{D}$ of problem-answer pairs $(x,a)$, which we call examples, a natural approach to maximize the average (over $\mathcal{D}$) reward is stochastic gradient ascent. At each iteration $t$, we update $\theta_{t+1}=\theta_{t}+\eta\cdot G(\theta_{t};\mathcal{D})$ with learning rate $\eta$ and gradient
$G(\theta;\mathcal{D})$ $=\widehat{\mathbb{E}}_{(x,a)\sim\mathcal{D}}[\nabla_{\theta}\rho_{\theta}(x,a)]$
$=\widehat{\mathbb{E}}_{(x,a)\sim\mathcal{D}}\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}\left[r_{0/1}(y,a)\cdot\nabla_{\theta}\log\pi_{\theta}(y|x)\right]\,$, (4)
where $\widehat{\cdot}$ denotes empirical averages and we have applied the log-derivative trick:
$\nabla_{\theta}\pi_{\theta}(y|x)=\pi_{\theta}(y|x)\cdot\nabla_{\theta}\log\pi_{\theta}(y|x)\,.$
REINFORCE-style algorithms provide different approximations of the per-example expected gradient $G_{0/1}(\theta;(x,a))=\nabla_{\theta}\rho_{\theta}(x,a)$.
REINFORCE.
The classical REINFORCE algorithm [Wil92] approximates the expectation over model outputs in (4) using the empirical average over $N$ sampled responses. This yields the per-example gradient update:
$\widehat{G}_{\text{REINFORCE}}(\theta;(x,a))=\frac{1}{N}\sum_{i\in[N]}r_{i}\cdot\nabla_{i}=\widehat{\rho}\cdot\widehat{\nabla}_{+}\,$. (5)
While simple, this estimator suffers from high variance and intuitively provides poor updates since it entirely ignores wrong responses.
RLOO.
REINFORCE with Leave-One-Out (RLOO) [KHW19, Ahm+24] reduces variance by replacing rewards $r_{i}$ with advantage scores that subtract a leave-one-out baseline:
$A_{i}=r_{i}-\frac{1}{N-1}\sum_{j\neq i}r_{j}\,.$
This also introduces learning signal from wrong responses. Since $r_{i}\in\{0,1\}$, the advantage scores depend only on whether response $y_{i}$ is correct or wrong:
$A^{+}=1-\frac{N\widehat{\rho}-1}{N-1}=\frac{N(1-\widehat{\rho})}{N-1}\qquad\text{and}\qquad A^{-}=-\frac{N\widehat{\rho}}{N-1}\,.$ (6)
Ignoring the common factor N/(N1)N/(N-1) across all examples, the RLOO gradient simplifies to:
[…]
This is an unbiased estimator of $\nabla_{\theta}\rho_{\theta}(x,a)$ with reduced variance compared to REINFORCE.
GRPO.
Group Relative Policy Optimization (GRPO) [Sha+24] normalizes advantages by their empirical standard deviation:
$A_{i}=\frac{r_{i}-\widehat{\rho}}{\sqrt{\widehat{\rho}(1-\widehat{\rho})}}\,$.
This yields correct/wrong advantage scores: [^1] Note that GRPO directly uses mean estimates across all responses (rather than leave-one-out). More importantly, it normalizes by standard deviation, making the gradient a biased estimator of $\nabla_{\theta}\rho_{\theta}(x,a)$.
$A^{+}=\frac{1-\widehat{\rho}}{\sqrt{\widehat{\rho}(1-\widehat{\rho})}}=\sqrt{\frac{1-\widehat{\rho}}{\widehat{\rho}}}\qquad\text{and}\qquad A^{-}=\frac{-\widehat{\rho}}{\sqrt{\widehat{\rho}(1-\widehat{\rho})}}=-\sqrt{\frac{\widehat{\rho}}{1-\widehat{\rho}}}\,$. (8)
Ignoring the clipping operation, the GRPO gradient is effectively [Den+25]:
$\widehat{G}_{\text{GRPO}}(\theta;(x,a))=\sqrt{\widehat{\rho}\cdot(1-\widehat{\rho})}\cdot\left$$\widehat{\nabla}_{+}-\widehat{\nabla}_{-}\right$]\,. (9)
[…]
3 Pass@K via Direct Differentiation
In analogy to the 0/1 case, the natural approach to optimize the expected Pass@K reward is stochastic gradient ascent with gradient
$G_{\text{\tiny Pass@K}}(\theta)=\widehat{\mathbb{E}}_{(x,a)\sim\mathcal{D}}[\nabla_{\theta}\rho_{K,\theta}(x,a)]\,.$
To compute the gradient, we rewrite the per-example Pass@K reward in terms of the per-example 0/1 reward. Using the IID nature of generated responses:
$\rho_{K,\theta}(x,a)$ $=\Pr_{\{y_{i}\}_{i\in[K]}\sim\pi_{\theta}(\cdot|x)}\left(\exists i\in[K]:y_{i}\text{ is correct}\right)=1-\prod_{i\in[K]}\Pr_{y_{i}\sim\pi_{\theta}(\cdot|x)}\left(y_{i}\text{ is \emph{not} correct}\right)$
$=1-(1-\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}r_{0/1}(y,a))^{K}\,.$ (10)
Following [Mah+25], we apply the log-derivative trick:
$\nabla_{\theta}\rho_{K,\theta}(x,a)$ $=-\nabla_{\theta}(1-\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}r_{0/1}(y,a))^{K}$
$=K(1-\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}r_{0/1}(y,a))^{K-1}\cdot\nabla_{\theta}\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}r_{0/1}(y,a)$
$=K(1-\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}r_{0/1}(y,a))^{K-1}\cdot\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}\left[r_{0/1}(y,a)\nabla_{\theta}\log\pi_{\theta}(y|x)\right]\,.$
Recognizing that the term in parentheses equals the per-example Fail@(K-1)=1=1-Pass@(K-1) reward, we obtain the final expression for the per-example (population) gradient:
$G_{\text{\tiny Pass@K}}(\theta;(x,a))$ $:=K\cdot\underbrace{\left(1-\rho_{K-1,\theta}(x,a)\right)}_{\text{Fail@(K-1)}}\,\cdot\,\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}\left[r_{0/1}(y,a)\nabla_{\theta}\log\pi_{\theta}(y|x)\right]\,.$ (11)
This is a reweighting of the respective gradient for the 0/1 reward. For $K=1$, Pass@1 equals the 0/1 reward and the gradient updates are indeed equivalent.
REINFORCEK.
With access to a finite number of responses $y_{1},\ldots,y_{N}$ per example, we need to approximate $G_{\text{\tiny Pass@K}}(\theta;(x,a))$ with an empirical estimate.
To arrive at an unbiased estimator, we first require an unbiased estimator for the Fail@(K-1) term in Eq. (11). The following is an unbiased estimator of $\rho_{K,\theta}(x,a)$ [Che+21]:
$\widehat{\rho}_{K}:=\widehat{\rho}_{K,\theta}(x,a)=1-{\binom{N^{-}}{K}}\Big/{\binom{N}{K}}\,$, (12)
representing the probability of drawing $K$ responses without replacement from $N$ total responses such that at least one is correct.
Having unbiased estimators for both terms in Eq. (11), we combine them via the leave-one-out trick to form an unbiased estimator of their product:
$\widehat{G}_{\text{REINFORCE}_{K}}(\theta;(x,a))=\frac{K}{N}\sum_{i\in[N]}(1-\widehat{\rho}_{K-1}^{\,\text{loo},i})\cdot r_{i}\cdot\nabla_{i},$ (13)
where $1-\widehat{\rho}_{K-1}^{\,\text{loo},i}$ is the leave-one-out unbiased estimator of Fail@(K-1) excluding the $i$-th response. For binary rewards, simple algebra (deferred to Appendix B.1) shows that
$1-\widehat{\rho}_{K-1}^{\,\text{loo},i}=\begin{cases}\widehat{f}_{K-1}^{+}:=(1-\widehat{\rho}_{K})\big/(1-\widehat{\rho}-\frac{K-1}{N})&\text{if }y_{i}\text{ is correct}\\ \widehat{f}_{K-1}^{-}:=(1-\widehat{\rho}_{K})\big/(1-\widehat{\rho})&\text{if }y_{i}\text{ is wrong}\end{cases}\,.$ (14)
The weights $\widehat{f}_{K-1}^{+}$ and $\widehat{f}_{K-1}^{-}$ are leave-one-out empirical estimates of Fail@(K-1), both expressed in terms of the empirical Pass@K reward $\widehat{\rho}_{K}$.
With these, the REINFORCEK update takes the following final form:
$\widehat{G}_{\text{REINFORCE}_{K}}(\theta;(x,a))$ $=\widehat{f}_{K-1}^{+}\cdot\widehat{\rho}\cdot\widehat{\nabla}_{+}=\widehat{f}_{K-1}^{+}\cdot\widehat{G}_{\text{REINFORCE}}(\theta;(x,a))\,.$ (15)
This is a reweighting of the vanilla REINFORCE gradient update [^2] We have dropped a constant $K$ factor across all examples, which can be absorbed into the learning rate. by $\widehat{f}_{K-1}^{+}$, the empirical Fail@(K-1) of the other N1N-1 samples: it amplifies gradients of examples with rare correct responses (large $N^{-}$) and suppresses gradients of redundant examples (small $N^{-}$). When $N^{-}<K-1$, the weight is 0 since Pass@K is already guaranteed.
As in the 0/1 case, REINFORCE forms an unbiased but high-variance estimator. We can reduce variance by substituting rewards $r_{i}$ in Eq. (13) with appropriately computed advantages. This way, we also introduce negative gradients.
RLOOK.
Using the RLOO advantage scores from Eq. (6) in place of $r_{i}$ in Eq. (13), the RLOO algorithm for Pass@K optimization has gradient (proportional to, ignoring constants N/(N1)N/(N-1) and $K$):
[…]
Note the slightly asymmetric weighting for correct versus wrong responses due to Pass@K maximization. This is an unbiased estimator of $G_{\text{\tiny Pass@K}}(\theta;(x,a))$ with lower variance compared to $\widehat{G}_{\text{REINFORCE}_{K}}(\theta;(x,a))$.
GRPOK.
Analogously, we obtain a GRPO update for Pass@K optimization by using the GRPO advantage scores from Eq. (8):
$\widehat{G}_{\text{GRPO}_{K}}(\theta;(x,a))=\widehat{f}_{K-1}^{+}\cdot\sqrt{\widehat{\rho}(1-\widehat{\rho})}\cdot\widehat{\nabla}_{+}-\widehat{f}_{K-1}^{-}\cdot\sqrt{\widehat{\rho}(1-\widehat{\rho})}\cdot\widehat{\nabla}_{-}\,$. (17)
4 Pass@K by Advantage Shaping
In Section 3, we saw that directly maximizing the Pass@K reward yields RLOOK and GRPOK variants that differ from their vanilla counterparts through asymmetric advantage reweighting. Specifically, the effective advantage score for Pass@K optimization is:
$A^{\pm}_{\text{\tiny Pass@K}}=\widehat{f}_{K-1}^{\pm}\cdot A^{\pm}\,$, (18)
where the weights $\widehat{f}_{K-1}^{\pm}$ differ for correct versus wrong responses (Eq. (14)), and $A^{\pm}$ are the vanilla advantage scores (e.g., Eq. (8) for GRPO).
[…]
$\widetilde{\text{GRPO}}_{K}$ via advantage shaping.
We now review the GRPO variant proposed by [Che+25], which we denote $\widetilde{\text{GRPO}}_{K}$, and which directly shapes GRPO’s advantages to favor Pass@K improvements. We refer readers to their nicely-written paper for details of the design rationale. Here, we begin from their proposed formulas for the modified GRPO advantages, which we denote $\widetilde{A}_{\text{\tiny Pass@K}}^{\pm}$. [Che+25, App. B3 ] propose the following scores:
$\widetilde{A}_{\text{\tiny Pass@K}}^{+}=\sqrt{\frac{1-\widehat{\rho}_{K}}{\widehat{\rho}_{K}}}\,\qquad\text{and}\qquad\widetilde{A}_{\text{\tiny Pass@K}}^{-}=\frac{1-\widehat{\rho}_{K}-\widehat{f}_{K-1}^{-}}{\sqrt{\widehat{\rho}_{K}(1-\widehat{\rho}_{K})}}\,$, (19)
expressed using our notation for the empirical Pass@K reward $\widehat{\rho}_{K}$ (Eq. (12)) and the leave-one-wrong-out Fail@(K-1) estimator $\widehat{f}_{K-1}^{-}$ (Eq. (14)).
As noted in [Den+25a], this can be simplified: Recall from Eq. (14) that $1-\widehat{\rho}_{K}=(1-\widehat{\rho})\cdot\widehat{f}_{K-1}^{-}$, which reveals $\widetilde{A}_{\text{\tiny Pass@K}}^{-}=-\frac{\widehat{\rho}}{1-\widehat{\rho}}\cdot\widetilde{A}_{\text{\tiny Pass@K}}^{+}$. Thus, in terms of vanilla GRPO advantage scores:
$\widetilde{A}_{\text{\tiny Pass@K}}^{\pm}=\widetilde{\omega}_{K}\cdot A^{\pm},\qquad\text{where }\,\,\,\,\,\widetilde{\omega}_{K}:=\sqrt{\frac{1-\widehat{\rho}_{K}}{\widehat{\rho}_{K}}}\cdot\sqrt{\frac{\widehat{\rho}}{1-\widehat{\rho}}}\,$. (20)
Overall, the per-example gradient of $\widetilde{\text{GRPO}}_{K}$ is:
$\widehat{G}_{{\text{$\widetilde{\text{GRPO}}_{K}$}}}(\theta;(x,a))=\widetilde{\omega}_{K}\cdot\widehat{G}_{\text{GRPO}}(\theta;(x,a))=\sqrt{\frac{1-\widehat{\rho}_{K}}{\widehat{\rho}_{K}}}\cdot\widehat{\rho}\cdot\left$$\widehat{\nabla}_{+}-\widehat{\nabla}_{-}\right$]\,. (21)
Comparing the two approaches.
Both methods reweight vanilla advantages and have a similar effect: compared to the vanilla methods of Sec. 2 that optimize 0/1 reward, they aggressively downweight gradient contributions from examples with medium-to-high empirical 0/1 rewards, thereby favoring more difficult examples. Yet, they differ structurally: The advantage-shaping reweighting in Eq. (20) is symmetric (same multiplier $\widetilde{\omega}_{K}$ for correct and wrong responses), whereas Eq. (18) uses asymmetric weights ($\widehat{f}_{K-1}^{+}\neq\widehat{f}_{K-1}^{-}$). Moreover, direct optimization more aggressively dampens contributions when the empirical reward is high (and amplifies more when low). Figure 1 visualizes the effective weights each algorithm applies to correct vs. wrong gradients.
5 Bridging the Views
5.1 Reverse-Engineering: Surrogate Reward from Advantage Shaping
In Section 4, we interpreted both Pass@K methods as rescalings of vanilla GRPO advantages: GRPOK via REINFORCE-style Pass@K reward optimization (Section 3) and $\widetilde{\text{GRPO}}_{K}$ via advantage-shaping. Both employ reweightings that account for empirical Pass@K performance. Here, we tighten the connection by showing that the advantage-shaping method also admits an interpretation as direct optimization of a surrogate Pass@K reward.
Claim 1 (Reverse-engineering $\widetilde{\text{GRPO}}_{K}$).
Recall $\widehat{\rho}:=\widehat{\rho}_{\theta}(x,a)$ denotes the per-example empirical 0/1 reward and $\widehat{\rho}_{K}:=\widehat{\rho}_{K,\theta}(x,a)$ denotes the per-example empirical estimate of the Pass@K reward (Eq. (12)). The method of [Che+25], $\widetilde{\text{GRPO}}_{K}$, rescales the vanilla GRPO empirical gradient yielding updates as shown in Eq. (21). For sufficiently large sample size $N\gg K$, this corresponds to direct maximization of the per-example surrogate reward
$\rho^{\prime}_{K,\theta}(x,a)\;=\;\frac{2}{K}\arcsin\!\Big(\sqrt{\rho_{K,\theta}(x,a)}\,\Big)\,$, (22)
where $\rho_{K,\theta}(x,a):=\mathbb{E}_{\{y_{i}\}_{i\in[K]}\sim\pi_{\theta}(\cdot|x)}\,r_{\text{\tiny Pass@K}}(\{y_{i}\}_{i\in[K]},a)$ is the expected Pass@K reward per example.
Thus, the advantage-shaping approach of [Che+25] implicitly optimizes a smooth, differentiable transformation of the Pass@K reward, providing further justification for the validity of the different approach by which the authors arrive at it. Note that the surrogate reward is maximized when the Pass@K reward equals 1, and is actually strictly monotone, so both the original and surrogate objectives share the same optimal solution, though the arcsin surrogate reward leads to a different optimization path.
Connection to variance stabilizing transforms.
The arcsin transform that emerges is noteworthy: it is a variance-stabilizing transformation (VST) for the binomial distribution. Specifically, for $\mathcal{X}\sim\text{Bin}(M,p)$, the arcsin transform satisfies $\text{Var}(\sqrt{M}\cdot\arcsin(\sqrt{\mathcal{X}/M}))\approx 1/4$ independent of $p$ [Ans48]. This connection is not coincidental: [Che+25, Sec. 2.2 ] arrived at their method through arguments using batched sampling, where $N$ responses are partitioned into $N/K$ groups of size $K$. In this setting, the number of successful groups follows a binomial distribution with $M=N/K$ trials, and the arcsin transform would stabilize the variance of this estimate. Formalizing the connection between VSTs and optimization stability when applying empirical gradients to transformed rewards is open direction for future work.
5.2 Forward-Engineering: Advantage Shaping from Surrogate Reward
Having shown that the $\widetilde{\text{GRPO}}_{K}$ method can be reverse-engineered as optimizing an implicit population-level surrogate reward, we now address the forward direction:
How does one arrive at $\widetilde{\text{GRPO}}_{K}$ starting from the arcsin-transform surrogate reward?
Claim 2 ($\widetilde{\text{GRPO}}_{K}$=RLOO on surrogate reward).
For $N\gg K$, the $\widetilde{\text{GRPO}}_{K}$ policy-gradient update by [Che+25] is equivalent to an RLOO-style policy gradient update to the surrogate per-example reward $\frac{2}{K}\arcsin\!\big(\sqrt{\rho_{K,\theta}(x,a)}\,\big)\,$, from Eq. (22).
To see this, we work as in Sec. 3, but now with the arcsin-transformed objective. We proceed in three steps: (i) By direct differentiation, the population per-example gradient is
$\frac{2}{K}\nabla_{\theta}\arcsin\!\big(\sqrt{\rho_{K,\theta}}\big)$ $=\frac{1}{K}\frac{1}{\sqrt{\rho_{K,\theta}(1-\rho_{K,\theta})}}\nabla_{\theta}\rho_{K,\theta}\,=\frac{1}{\sqrt{\rho_{K,\theta}(1-\rho_{K,\theta})}}(1-\rho_{\theta})^{K-1}\nabla_{\theta}\rho_{\theta}$
$=\frac{1}{\sqrt{\rho_{K,\theta}(1-\rho_{K,\theta})}}\cdot\frac{1-\rho_{K,\theta}}{1-\rho_{\theta}}\cdot\nabla_{\theta}\rho_{\theta},$ (23)
To obtain an empirical version of this update, we (ii) approximate the multiplicative term of the gradient by substituting $\rho_{K,\theta}$ and $\rho_{\theta}$ with their empirical counterparts $\widehat{\rho}_{K}$ and $\widehat{\rho}$, and (iii) approximate the expected 0/1 reward gradient term with the RLOO gradient update from Eq. (16) (recall Eq. (1)). The update then becomes:
$\frac{1}{\sqrt{\widehat{\rho}_{K}(1-\widehat{\rho}_{K})}}\cdot\frac{1-\widehat{\rho}_{K}}{1-\widehat{\rho}}\cdot\widehat{\rho}(1-\widehat{\rho})\cdot\left[\widehat{\nabla}_{+}-\widehat{\nabla}_{-}\right]=\sqrt{\frac{1-\widehat{\rho}_{K}}{\widehat{\rho}_{K}}}\cdot\widehat{\rho}\cdot\left$$\widehat{\nabla}_{+}-\widehat{\nabla}_{-}\right$]\,,
and we recognize $\widehat{G}_{{\text{$\widetilde{\text{GRPO}}_{K}$}}}(\theta;(x,a))$ in the final expression (see Eq. (21)).
This completes a bidirectional link between surrogate optimization and advantage shaping: $\widetilde{\text{GRPO}}_{K}$ can both be reverse-engineered as optimizing an arcsin-transformed Pass@K reward and forward-constructed by applying RLOO to this surrogate.
Alternative surrogate rewards.
This connection invites exploration of alternative variance-stabilizing transformations. The arcsin transform is a VST for the binomial distribution, which aligns with the batched sampling scheme of [Che+25, Sec. 2.2 ]. However, the standard Pass@K estimator (Eq. (12)) is combinatorial and does not strictly follow a binomial distribution. Deriving a VST tailored to this estimator’s distribution, then applying the forward-engineering recipe from Claim 2, could yield a more stable optimization algorithm.
Related to this, the analysis above assumes $N$ grows large while $K$ remains constant. Designing surrogate rewards that account for joint scaling of $N$ and $K$ is an interesting direction, particularly for the practically relevant small-$N$ regime where computational budgets are limited.
Biased versus unbiased estimation.
The derivation in Claim 2 differs from the RLOOK construction in Sec. 3 in that we do not insist on maintaining unbiasedness of the gradient estimator. For example, $\frac{1-\widehat{\rho}_{K}}{1-\widehat{\rho}}$ is a biased estimator of the multiplicative term in Eq. (23); even replacing $\widehat{\rho}_{K}$ with a leave-one-out estimate would not remove this bias since the product of two unbiased estimators is generally biased. This is perhaps to be expected: even the strong vanilla GRPO baseline does not implement an unbiased empirical estimate of the gradient.
In fact, applying the logic of Claim 2 to the $K=1$ case provides an alternative interpretation to GRPO’s normalization by the reward’s standard deviation.
Corollary 1 (GRPO as Surrogate Reward Optimization).
For $N\gg 1$, the (vanilla) GRPO policy-gradient update (Eq. (9)) by [Sha+24] is equivalent to an RLOO-style policy gradient update for the surrogate per-example reward $2\arcsin\!\Big(\sqrt{\rho_{\theta}(x,a)}\,\Big)\,$, where $\rho_{\theta}(x,a):=\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}\,r_{0/1}(y,a)$ is the expected 0/1 reward per example.
[…]
6 A Regularized Surrogate Rewards Percpective
Both Pass@K routes examined in Secs. 3 and 4 tilt the learning signal toward the hard cases. [Che+25] argue this is aligned with the exploit/explore roles of the two metrics. Maximizing the 0/1 reward emphasizes exploitation: making individual responses more often correct, even on already-easy examples. In contrast, Pass@K encourages exploration: for a given example, only one of $K$ responses must be correct, so once an example is “solved” in the Pass@K sense, further pushing its 0/1 reward is less valuable than allocating gradient signal to unsolves examples where more search could pay off. Ideally, we want a controlled balance: retain sufficient signal on solved examples while injecting additional search signal on unsolved ones. To achieve this, they propose mixing Pass@K-shaped and vanilla advantages.
Using reverse-engineering, we arrive at a complementary interpretation of their method as regularized surrogate reward maximization. This perspective decomposes the RL objective into a data-fitting term (the transformation of the mean reward that the GRPO baseline optimizes) and a regularizer (an uncertainty term such as the reward’s standard deviation). While the data-fitting term can be interpreted as exploitation and the regularizer as exploration, the regularization viewpoint itself provides a possibly more elementary justification for practical advantage-shaping heuristic-modifications of GRPO.
6.1 Combined 0/1 and Pass@K Advantage Shaping
Having observed that Pass@K optimization can aggressively favor difficult examples (small $\widehat{\rho}$) while potentially zeroing out useful signal on easier ones (medium/large $\widehat{\rho}$), [Che+25] propose a simple blend between optimizing the $0/1$ reward and optimizing Pass@K. They combine the Pass@K advantage with the vanilla advantage using the empirical 0/1 reward $\widehat{\rho}$ as a mixing weight. In our notation:
$\widetilde{A}_{\text{mix}}^{\pm}\;=\;\widehat{\rho}\,\widetilde{A}_{\text{\tiny Pass@K}}^{\pm}\;+\;(1-\widehat{\rho})\,A^{\pm}\;=\;\bigl(1-\widehat{\rho}\,+\widehat{\rho}\cdot\widetilde{\omega}_{K}\bigr)\,A^{\pm},$ (24)
where we used Eq. (20) to write $\widetilde{A}_{\text{\tiny Pass@K}}^{\pm}=\widetilde{\omega}_{K}\,A^{\pm}$.
The same idea can be instantiated for the direct Pass@K optimization view presented in Sec. 3. Since $A_{\text{\tiny Pass@K}}^{\pm}=\widehat{f}_{K-1}^{\pm}\,A^{\pm}$ by (18), an analogous combination yields
$A_{\text{mix}}^{\pm}\;=\;\widehat{\rho}\,A_{\text{\tiny Pass@K}}^{\pm}\;+\;(1-\widehat{\rho})\,A^{\pm}\;=\;\bigl(1-\widehat{\rho}\,+\widehat{\rho}\cdot\widehat{f}_{K-1}^{\pm}\bigr)\,A^{\pm}.$ (25)
Figure 2 visualizes the resulting effective gradient weights (i.e., $\widehat{\rho}A^{+}_{\text{mix}}$ for correct and $(1-\widehat{\rho})\lvert A^{-}_{\text{mix}}\rvert$ for wrong responses). First observe this blending avoids the overdamping of mid/large-$\widehat{\rho}$ training signal that can occur under pure Pass@K optimization (whether direct or via advantage shaping). Interestingly, the two combinations become numerically very close as $K$ grows. Moreover, except when $K$ is small relative to $N$, both essentially collapse to the lower envelope $(1-\widehat{\rho})A^{\pm}$. That is, they essentially no longer exploit an explicit Pass@K signal.
Figure 2: Combined Pass@K effective gradient scores for $N=16$ and $K\in\{2,4,6,8,10\}$. Green uses Eq. (25); Blue uses Eq. (24); the dotted gray line is the baseline $(1-\widehat{\rho})A^{\pm}$ (which we call skew-R). For $K=2,4$ the two combinations differ slightly at mid/high $\widehat{\rho}$ (blue is modestly larger); for $K\geq 6$ they nearly coincide across $\widehat{\rho}$. This is because, for most $\widehat{\rho}$, the Pass@K scalers vanish for $\widehat{\rho}<1-K/N$, reducing both combinations to the same envelope $(1-\widehat{\rho})\sqrt{\widehat{\rho}(1-\widehat{\rho})}$.
6.2 Reward Regularization Interpretation of skew-R
What reward does the combined 0/1 and Pass@K advantage shaping actually optimize?
Because Sec. 4 connects advantage shaping to direct Pass@K optimization, one might guess it maximizes a convex combination of the $0/1$ and Pass@K rewards. The catch is that in (24) the mixing weight $\widehat{\rho}$ is adaptive to the instance “difficulty” (as measured by $\widehat{\rho}$), so the combination is not a fixed convex blend at the objective level.
To simplify reverse engineering and obtain a more interpretable surrogate reward, we leverage the empirical observation from the previous subsection: except when $K$ is small relative to $N$, both combined schemes $\widetilde{A}^{\pm}_{\text{mix}}$ and $A^{\pm}_{\text{mix}}$ are numerically close in terms of effective weights to the simpler right-skew shaping
$A^{\pm}_{\text{skew-R}}\;:=\;(1-\widehat{\rho})\,A^{\pm},$
which no longer uses an explicit Pass@K signal but reproduces a shape that approximates the mixture with Pass@K.
Claim 3 (Skew-R== regularized surrogate reward maximization).
[…]
$\widehat{G}_{\text{skew-R}}(\theta;(x,a))\;=\;(1-\widehat{\rho})\cdot\widehat{G}_{\text{GRPO}}(\theta;(x,a))\,.$ (26)
When $N\gg 1$, this corresponds to maximization of the per-example reward
$\arcsin\!\Big(\sqrt{\rho}\,\Big)\;+\;\sqrt{\rho\,(1-\rho)}\,$, (27)
where $\rho:=\rho_{\theta}(x,a):=\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}[r_{0/1}(y,a)]$ is the (population) 0/1 reward per example.
Reward-regularization perspective.
Vanilla REINFORCE and RLOO maximize the expected 0/1 reward $\rho$ (pure exploitation). As we saw in Corollary 1, vanilla GRPO maximizes a strictly monotone transform of the mean, $\arcsin(\sqrt{\rho})$, so it has the same maximizers as $\rho$ and likewise corresponds to exploitation. In contrast, the skew-R reward can be interpreted as stochastic gradient ascent on a regularized surrogate reward that combines a data-fitting term and a regularizer. The data-fitting term incentivizes finding a policy that maximizes the reward over the training examples $(x,a)\in\mathcal{D}$. In contrast, the regularizer alone is maximized at $\rho=1/2$ where the binary correct and wrong policy outcomes are equally probable. An intuitive explanation of this term, aligning with the regularization effect in classical supervised learning, is that it incentivizes the model to not “overfit” on the training examples, maintaining diversity in wrong response that may explore different solution paths generalizing to unseen examples. Formalizing such a notion of overfitting and its connection to regularization could provide deeper insights into RLVR learning tradeoffs.
We note that the argmax of the combined reward in Eq. (27) is still at $\rho=1$, reassuring here that we incentivize ultimately finding parameters that maximize the average reward. However, mirroring lessons from supervised fine-tuning, this might not always be the best strategy for generalization, and stronger regularization with an explicit parameter controlling the strength could be beneficial (see next subsection).
Implications for the “Prioritized Sampling” strategy of Kimi 1.5.
An interesting observation potentially supporting the practical value of skew-R: Kimi 1.5 [Tea+25] performs priority sampling in their RL experiments, reweighting each example by $1-\widehat{\rho}$ to allow harder examples to appear more frequently. This corresponds to the advantage shaping of skew-R (Eq. (27)) applied at the example level rather than within gradient computation. Thus, our interpretation as regularized reward maximization provides a direct justification for their empirically motivated practical choice.
Consistency of interpretations.
In the large-sample limit $N\gg 1$, skew-R is equivalent to GRPOK=2. In Appendix C.2, we elaborate on this comparison with emphasis on interpreting the reverse-engineered surrogate reward. We show that GRPOK=2 effectively optimizes an incomplete-beta-function transformation of the Pass@2 ($\rho_{2}$) reward, ${\mathcal{B}}\!\big(1-(1-\rho_{2})^{1/2};\tfrac{1}{2},\tfrac{3}{2}\big),$ and that this objective is algebraically identical to the skew-R surrogate in Eq. (27). Thus, in the population limit, both methods are in fact maximizing the same scalar objective.
This validates our surrogate-reward reverse-engineering framework, but it also highlights a subtle interpretive question about what constitutes “regularization.” The interpretation depends on which performance metric we take as primary:
  • Pass@K-centric lens. If Pass@K ($\rho_{K}$) is treated as the true performance metric, then GRPOK should be viewed as directly maximizing a smooth, monotone surrogate of $\rho_{K}=1-(1-\rho)^{K}$. In this view there is no extra penalty term: we are simply doing surrogate maximization of the evaluation metric itself.
[…]
Both lenses are mathematically consistent; the choice is purely about framing of the learning goal.
Figure 3: Comparison of (regularized) surrogate objectives (GRPO, Skew-R, and Entropy-Augmented). (a) Effective Gradient Weight: Empirical gradient’s scaling factor as a function of the empirical reward $\hat{\rho}$. The vanilla GRPO (blue, dashed) is symmetric. Skew-R (green) and the Entropy-augmented methods (red, purple) are asymmetric, suppressing gradients for high $\hat{\rho}$. Skew-R gradients are multiplied here by $2$, so all four curves have the same area under the curve. (b) Surrogate Reward Value: Corresponding surrogate reward functions, normalized so that they evaluate to 11 at $\rho=1$.
6.3 An Example of Reward-Level Regularization
The analysis of skew-R revealed that a simple gradient scaling, $(1-\widehat{\rho})\cdot\widehat{G}_{\text{GRPO}}$, of vanilla GRPO admits a population interpretation as policy gradient optimization with a regularizer term added on the ”good” baseline surrogate $\arcsin(\sqrt{\rho_{\theta}})$ of the original GRPO.
What if we use a different regularizer? What is the effect on the shape of the advantage scores?
To illustrate the use of reward regularization, we show how to optimize a surrogate reward that combines the data-fitting term with an entropy regularizer $H(\rho)=-\rho\log(\rho)-(1-\rho)\log(1-\rho)$ of the binary 0/1 reward. Intuitively, similar to the std-regularization of skew-R, $H(\rho)$ favors examples where the model is still “undecided,” redistributing gradient effort toward those rather than repeatedly polishing ones it already gets right. The overall surrogate reward is
$2\,\underbrace{\arcsin(\sqrt{\rho})}_{\text{GRPO's baseline}}+\;\;\lambda\cdot\underbrace{H(\rho)}_{\text{Regularizer}}\,.$
Applying the forward-engineering recipe (differentiate the per-example regularized objective above, then replace population quantities by their empirical estimates $\hat{\rho}$ and plug in the usual low-variance GRPO estimator for $\nabla_{\theta}\rho$; see Appendix B.4) yields a tunable advantage-shaping rule that scales the vanilla GRPO gradient as follows:
$\widehat{G}_{\text{entropy}}={\left[1-\lambda\cdot\sqrt{\widehat{\rho}(1-\widehat{\rho})}\log\left(\frac{\widehat{\rho}}{1-\widehat{\rho}}\right)\right]}\cdot\,\widehat{G}_{\text{GRPO}}\,.$ (28)
[…]
Note that unlike Skew-R’s objective in Eq. (27), which is maximized at $\rho_{\theta}=1$, the entropy-augmented objective need not be globally maximized at $\rho=1$ for all $\lambda$. For sufficiently small $\lambda\lesssim 1.5$, the maximizer of $J_{\lambda}$ remains at $\rho=1$, so we are still ultimately pushing toward perfect success. For larger regularization strengths, the entropy term is strong enough that the overall maximizer shifts to some $\rho_{\lambda}<1$, providing a mechanism to penalize perfect fitting of the training data, a stronger form of regularization.
7 Practical Considerations
7.1 Biased vs Unbiased Scalings
In Sec. 3 we applied the log-derivative trick directly to the Pass@K objective, yielding the expected gradient $G_{\text{\tiny Pass@K}}(\theta;(x,a))$ in Eq. (11). We then derived Pass@K variants of REINFORCE by using a leave-one-out construction to obtain unbiased estimates of the two components in (11): $\mathrm{Fail}@{(K-1)}$ and $G_{0/1}(\theta;(x,a))$, resulting in the unbiased empirical gradient $\widehat{G}_{\text{RLOO}_{K}}$ in Eq. (16).
In contrast, [Mah+25] apply a biased empirical gradient that estimates the $\mathrm{Fail}@{(K-1)}$ factor as $(1-\widehat{\rho})^{K-1}$. The corresponding advantage-score scaling is (up to constant multiplicative factor $K$)
$(1-\widehat{\rho})^{K-1}\cdot A^{\pm}\,.$ (29)
Unlike the estimators in Sec. 3, this remains biased even if the advantage-scores $A^{\pm}$ correspond to an unbiased method (e.g., RLOO), because $(1-\widehat{\rho})^{K-1}$ is itself a biased estimator of $\mathrm{Fail}@{(K-1)}$. Replacing it by a leave-one-out version $(1-\widehat{\rho}^{\mathrm{loo}})^{K-1}$ does not remove this bias.
Note that such biased scaling is also present in advantage-shapes resulting from our forward-engineering of surrogate rewards in Sec. 5. Even our reverse-engineering operates at the population level. Thus, here, it would tell us that both GRPOK (which applies unbiased reweightings to vanilla GRPO) and the method of [Mah+25], both implicitly maximize the same surrogate reward (see Sec. C.1). In the practical regime of few samples though the two methods could lead to different behaviors. While unbiasedness is intuitively a good property, we discuss below regimes where the biased variant can be particularly attractive. It is worth noting that for $K=2$ the two variants are identical as they coincide with skew-R, which as we discussed in Sec. 5 is also effectively used by [Tea+25].
Figure 4: Comparison of absolute effective gradient weights (log scale) for Pass@K training with $N=16$. Each panel shows a different $K$ and plots the magnitude applied to correct (solid) and wrong (dashed) responses under three schemes: GRPOK (Eq. (18); green), $\widetilde{\text{GRPO}}_{K}$ (Eq. (20); blue), and the biased scaler (Eq. (29); red). For small $\hat{\rho}$, GRPOK assigns slightly larger weights than the biased scaler. As $\hat{\rho}$ increases, GRPOK enforces a hard zero once $\hat{\rho}>1-\frac{K-1}{N}$, whereas the biased scaler decays smoothly. $\widetilde{\text{GRPO}}_{K}$ applies a symmetric, example-level scale and often yields the largest mid-$\hat{\rho}$ weights until $\hat{\rho}_{K}$ saturates. Only the GRPOK exhibits asymmetric scaling between correct and wrong responses.
When a biased reweighting factor can help.
There are practical regimes where biased scalings are intuitively attractive. Specifically, for Pass@K optimization this includes (see also Fig. 4): (1) Compute-constrained $N\!\leq\!K$. The unbiased Pass@K estimator in Sec. 3 (and the grouping-based method of Sec. 4) is undefined when $K>N$, whereas $(1-\widehat{\rho})^{K-1}$ still yields a meaningful signal. (2) Small $N$ relative to $K$. When $N>K$ but not by much, the unbiased combinatorial weight becomes exactly zero whenever $N^{-}<K-1$, prematurely killing gradients; the smooth scaler $(1-\widehat{\rho})^{K-1}$ avoids this early saturation and provides a gentler shaping $A_{\text{\tiny Pass@K}}^{\pm}=(1-\widehat{\rho})^{K-1}A^{\pm}$ that maintains nonzero updates. (3) Very large $N$ relative to $K$. When $N\gg K$, $\widehat{f}_{K-1}^{+}\approx\widehat{f}_{K-1}^{-}\approx\bigl(N^{-}/(N-1)\bigr)^{K-1}\approx(1-\widehat{\rho})^{K-1}$, so the two estimators nearly coincide.
7.2 Normalization Factors
Up to the min-clip operation, the main difference of GRPO from classical RLOO is normalization with respect to the reward’s standard deviation (Sec. 2). The surrogate reward perspective (Cor. 1) provides an alternative interpretation of this scaling: at the level of the training objective, it maximizes a transformation of the 0/1 reward that stabilizes the variance of its empirical estimator. This interpretation is consistent for $\widetilde{\text{GRPO}}_{K}$, which as we discussed in Sec. 4 was originally designed through advantage shaping in a batched setting where Pass@K is evaluated on $N/K$ batches of the observed responses. More broadly, our forward-engineering of reward surrogates $F(\rho)$ uses RLOO empirical gradients as a robust estimate of the population reward gradient $\nabla_{\theta}\rho$ and scales them by $F^{\prime}(\rho)$ that generalizes the specific GRPO normalization via standard deviation.
On the empirical side, it is worth noting that there is ongoing debate about whether such variance-stabilizing normalizations are necessary for good performance. For example, in support of earlier works [Ahm+24], a recent study by [Kha+25] finds that RLOO is competitive to GRPO, with zero-mean advantages being the main driver behind its success [Xio+25]. This raises an interesting question: Could we theoretically characterize how properties of a surrogate $F$ (e.g., variance stabilization) relate to favorable optimization dynamics? Such an analysis might shed light on if, when and why these normalizations matter in practice.
7.3 Granular Advantage Shaping
[…]
One can imagine even finer-grained shaping. For example, [wenlong_THR] investigate token-level advantage shaping through the lens of simple one-step dynamics of vanilla expected reward maximization: token-level control incentivizes the model to preserve probability mass on unobserved alternative responses to avoid over-concentrating on observed ones as a means of balancing exploitation and exploration. The entropy-based modification of advantages by [Che+25a] is effectively another instance of more granular token-level control introduced to favor exploration. See also [Cui+25] for yet another token-level advantage shaping motivated by detailed empirical investigation of how entropy relates to downstream reasoning performance.
It would be interesting to apply the reverse-engineering perspective to interpret these token-level methods as regularized surrogate rewards. The finer granularity naturally suggests finer-grained regularization at the policy or even model-representation level. Though this would be more challenging to formalize, a common surrogate-reward perspective could provide an alternate lens to actual benefits brought by the various modifications.
7.4 Clipping
Our analysis omits the min-clip operation of GRPO [Sha+24], which is itself inherited from PPO [Sch+17]. This is accurate when GRPO operates in full online mode (single update per model generation), as originally done by [Sha+24]. However, if multiple updates are performed per generation, the GRPO gradient expression in Eq. (9) that our analysis uses is only a proxy. That said, all advantage-shaping methods studied here (e.g., GRPOK, skew-R, entropy-augmented) can be empirically applied by simply substituting the respective advantage scores into the original GRPO formulation with the min-clip operation (or follow-up variants, e.g., [Yu+25]).
8 Conclusion
Motivated by an effort to reconcile recent Pass@K policy gradient methods, we show that advantage shaping is (asymptotically) equivalent to RLOO-style optimization of surrogate rewards $F(\rho)$. This equivalence extends beyond Pass@K: for instance, “hard-example upweighting” heuristics correspond to adding reward-level regularization $\Omega(\rho)$ to the training objective. While $F(\rho)$ is maximized at $\rho=1$, acting as a data-fitting term, the regularizer $\Omega(\rho)$ provides an opposing force that can be interpreted as preventing overfitting, particularly to easy training examples.
Advantage shaping and surrogate reward design are two sides of the same coin. This equivalence mirrors classical supervised learning, where modifying the training objective is analogous to directly modifying the gradient. Both design approaches have proven effective, with a strong parallel in supervised classification for imbalanced data, e.g. [Lin+17, Men+20, YZC21, Kin+21]. At a conceptual level, understanding this equivalence is instructive: irrespective of the design choice—whether starting from advantage shapes or surrogate rewards—analyzing the counterpart can provide valuable insight. Looking forward, while direct parallels to supervised learning should be drawn cautiously, the (regularized) surrogate reward lens may offer a bridge between iterative optimization algorithms, like policy gradients, and the population performance metrics they ultimately aim to optimize.
Acknowledgements
This work was partially funded by the NSERC Discovery Grant No. 2021-03677 and the Alliance Grant ALLRP 581098-22.
Appendix A Auxiliary Technical Results
Lemma 1 (Relating $\widehat{f}_{K-1}^{+}$, $\widehat{f}_{K-1}^{-}$, and $\widehat{\rho}_{K}$).
With
$\widehat{\rho}_{K}\;=\;1-\frac{\binom{N^{-}}{K}}{\binom{N}{K}},\qquad\widehat{f}_{K-1}^{+}\;=\;\frac{\binom{N^{-}}{K-1}}{\binom{N-1}{K-1}},\qquad\widehat{f}_{K-1}^{-}\;=\;\frac{\binom{N^{-}-1}{K-1}}{\binom{N-1}{K-1}}$,
and the conventions $\binom{N}{k}=0$ when $N<k$ and $\binom{N}{0}=1$, we have
$\quad 1-\widehat{\rho}_{K}\;=\;(1-\widehat{\rho})\cdot\,\widehat{f}_{K-1}^{-}\;=\;(1-\widehat{\rho}-\frac{K-1}{N})\cdot\,\widehat{f}_{K-1}^{+}\quad$}.
Proof.
Use the standard identities $\binom{N}{k}=\frac{N}{k}\binom{N-1}{k-1}$ and $\binom{N}{K}=\frac{N}{K}\binom{N-1}{K-1}$. Then
$\frac{\binom{N^{-}}{K}}{\binom{N}{K}}=\frac{\frac{N^{-}}{K}\binom{N^{-}-1}{K-1}}{\frac{N}{K}\binom{N-1}{K-1}}=\frac{N^{-}}{N}\cdot\frac{\binom{N^{-}-1}{K-1}}{\binom{N-1}{K-1}}=\frac{N^{-}}{N}\,\widehat{f}_{K-1}^{-},$
which implies $\widehat{\rho}_{K}=1-\frac{N^{-}}{N}\,\widehat{f}_{K-1}^{-}$.
Alternatively, using $\binom{N-1}{k}=\frac{N-k}{N}\binom{N}{k}$ with $N=N^{-}$ and k=K1k=K-1,
$\widehat{f}_{K-1}^{-}=\frac{\binom{N^{-}-1}{K-1}}{\binom{N-1}{K-1}}=\frac{N^{-}-(K-1)}{N^{-}}\cdot\frac{\binom{N^{-}}{K-1}}{\binom{N-1}{K-1}}=\frac{N^{-}-(K-1)}{N^{-}}\,\widehat{f}_{K-1}^{+}.$
Substitute this into the previous display to get
$\frac{\binom{N^{-}}{K}}{\binom{N}{K}}=\frac{N^{-}}{N}\cdot\frac{N^{-}-(K-1)}{N^{-}}\,\widehat{f}_{K-1}^{+}=\frac{N^{-}-(K-1)}{N}\,\widehat{f}_{K-1}^{+},$
hence $\widehat{\rho}_{K}=1-\frac{N^{-}-(K-1)}{N}\,\widehat{f}_{K-1}^{+}$. The boundary conventions cover cases $K=1$ (both weights =1=1) and $M^{-}<K$ (ratio =0=0). ∎
Lemma 2 (Conditional-gradient identity).
Let $r:=r_{0/2}(y,a)$ and $\rho:=\rho_{\theta}(x,a):=\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}[r_{0/1}(y,a)]$. Define
$\Delta\;:=\;\mathbb{E}\!\left[\nabla_{\theta}\log\pi_{\theta}(y|x)\mid r=1\right]\;-\;\mathbb{E}\!\left[\nabla_{\theta}\log\pi_{\theta}(y|x)\mid r=0\right].$
Then
$\Delta\;=\;\frac{1}{\rho(1-\rho)}\,\nabla_{\theta}\rho.$
Proof.
By conditioning and the log-derivative trick,
$\mathbb{E}\!\left[\nabla_{\theta}\log\pi_{\theta}(y|x)\mid r=1\right]=\frac{\mathbb{E}[r\,\nabla_{\theta}\log\pi_{\theta}(y|x)]}{\rho}=\frac{\nabla_{\theta}\rho}{\rho}.$
Using $\mathbb{E}[\nabla_{\theta}\log\pi_{\theta}(y|x)]=0$,
$\mathbb{E}\!\left[\nabla_{\theta}\log\pi_{\theta}(y|x)\mid r=0\right]=\frac{\mathbb{E}[(1-r)\,\nabla_{\theta}\log\pi_{\theta}(y|x)]}{1-\rho}=-\,\frac{\nabla_{\theta}\rho}{1-\rho}.$
Subtracting gives $\Delta=\big(\tfrac{1}{\rho}+\tfrac{1}{1-\rho}\big)\nabla_{\theta}\rho=\frac{1}{\rho(1-\rho)}\,\nabla_{\theta}\rho.$
Appendix B Missing Proofs
B.1 Proof of Equation (14)
The leave-one-out unbiased estimator of Fail@(K-1) excluding response $i$ is:
$1-\widehat{\rho}_{K-1}^{\,\text{loo},i}=\frac{\binom{(N-1)(1-\widehat{\rho}^{\,\text{loo},i})}{K-1}}{\binom{N-1}{K-1}}\,$,
where the leave-one-out empirical 0/1 reward is
$\widehat{\rho}^{\,\text{loo},i}=\begin{cases}\frac{N^{+}-1}{N-1}&\text{if }y_{i}\text{ is correct}\\ \frac{N^{+}}{N-1}&\text{if }y_{i}\text{ is wrong}\end{cases}.$
For correct responses: $(N-1)(1-\widehat{\rho}^{\,\text{loo},i})=(N-1)-(N^{+}-1)=N^{-}$.
For wrong responses: $(N-1)(1-\widehat{\rho}^{\,\text{loo},i})=(N-1)-N^{+}=N^{-}-1$.
Apply now Lemma 1 to yield Eq. (14).
B.2 Proof of of Claim 1: Reverse-Engineering $\widetilde{\text{GRPO}}_{K}$
Recall the empirical update for the advantage-shaping method from Eq. (20):
$\widehat{G}_{\text{$\widetilde{\text{GRPO}}_{K}$}}$ $=\widetilde{\omega}_{K}\cdot\widehat{G}_{\text{GRPO}}=\sqrt{\frac{1-\widehat{\rho}_{K}}{\widehat{\rho}_{K}}}\sqrt{\frac{\widehat{\rho}}{1-\widehat{\rho}}}\cdot\widehat{G}_{\text{GRPO}}\,$
[…]
Replacing empirical $\hat{\cdot}$ quantities with their population counterparts, the population-level gradient (call it) $G_{\text{GRPO}_{K}}$ is:
$G_{\text{$\widetilde{\text{GRPO}}_{K}$}}=\left(\sqrt{\frac{1-\rho_{K}}{\rho_{K}}}\cdot\sqrt{\frac{\rho}{1-\rho}}\right)\sqrt{\rho(1-\rho)}\cdot\mathbb{E}_{y_{i}\sim\pi_{\theta}(\cdot|x)}\left[\widehat{\nabla}_{+}-\widehat{\nabla}_{-}\right$]\,.
The expectation in the last term becomes
$\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}\!\left[\nabla_{\theta}\log\pi_{\theta}(y|x)\mid r(y,a)=1\right]\;-\;\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}\!\left[\nabla_{\theta}\log\pi_{\theta}(y|x)\mid r(y,a)=0\right].$
Using Lemma 2, this equals $\nabla_{\theta}\rho/(\rho(1-\rho))$. Thus,
$G_{\text{$\widetilde{\text{GRPO}}_{K}$}}=\left(\sqrt{\frac{1-\rho_{K}}{\rho_{K}}}\cdot\sqrt{\frac{\rho}{1-\rho}}\right)\cdot\left(\frac{1}{\sqrt{\rho(1-\rho)}}\nabla_{\theta}\rho\right)=\left(\sqrt{\frac{1-\rho_{K}}{\rho_{K}}}\cdot\frac{1}{1-\rho}\right)\cdot\nabla_{\theta}\rho\,.$ (31)
We now relate $\rho$ to $\rho_{K}$. Recall from Eq. (10) that $\rho_{K}=1-(1-\rho)^{K}\implies 1-\rho=(1-\rho_{K})^{1/K}$. From this also $\nabla_{\theta}\rho_{K}=K(1-\rho)^{K-1}\nabla_{\theta}\rho=K(1-\rho_{K})^{1-1/K}\nabla_{\theta}\rho\,.$
Substitute these two relationships back into (31):
$G_{\text{$\widetilde{\text{GRPO}}_{K}$}}$ $=\left(\frac{1}{K}\cdot\sqrt{\frac{1-\rho_{K}}{\rho_{K}}}\cdot\frac{1}{(1-\rho_{K})^{1/K}(1-\rho_{K})^{1-1/K}}\right)\cdot\nabla_{\theta}\rho_{K}$
$=\left(\frac{1}{K\sqrt{\rho_{K}(1-\rho_{K})}}\right)\cdot\nabla_{\theta}\rho_{K}\,.$ (32)
This is now in the form $G_{\text{$\widetilde{\text{GRPO}}_{K}$}}=F(\rho_{K})\nabla_{\theta}\rho_{K}$. We integrate in $\rho_{K}$ to find the surrogate $F$:
$\int\frac{1}{\sqrt{\rho_{K}(1-\rho_{K})}}\,d\rho_{K}=2\arcsin(\sqrt{\rho_{K}})+C=2\arcsin(2\rho_{K}-1)+C^{\prime}\,$, (33)
where $C,C^{\prime}$ are constants. This completes the proof of the claim.
B.3 Proof of Claim 3: Reverse-engineering Skew-R
Recall the empirical GRPO update (no clipping) from Eq. (9):
$\widehat{G}_{\text{GRPO}}=\sqrt{\widehat{\rho}(1-\widehat{\rho})}\left$$\widehat{\nabla}_{+}-\widehat{\nabla}_{-}\right$].
We have seen in the proof of Claim 1 that, at the population level this update becomes
$G_{\text{GRPO}}\;=\;\frac{1}{\sqrt{\rho(1-\rho)}}\,\nabla_{\theta}\rho,$
Then, the skew-R update at the population level is
$G_{\text{skew-R}}\;=\;(1-\rho)\,G_{\text{GRPO}}\;=\;\sqrt{\frac{1-\rho}{\rho}}\;\nabla_{\theta}\rho.$
We can more conveniently express this as
$G_{\text{skew-R}}=\frac{1}{\sqrt{2\rho(1-\rho)}}\cdot\nabla_{\theta}\rho+\frac{1-2\rho}{\sqrt{2\rho(1-\rho)}}\cdot\nabla_{\theta}\rho\,.$
We know from Corollary 1 that the first term corresponds to a surrogate reward $\arcsin(\sqrt{\rho})$. Analogously, the second term corresponds to $\Omega(\rho)$ for $\Omega$ such that $\Omega^{\prime}(\rho)=\frac{1-2\rho}{\sqrt{2\rho(1-\rho)}}$. Integrating gives (up to a constant) $\Omega(\rho)=\sqrt{\rho(1-\rho)}$ completing the proof of the claim.
B.4 Reward-Level Entropy Regularization
We start by computing the population gradient of the per example entropy-surrogate reward $2\arcsin(\sqrt{\rho})+\lambda\cdot H(\rho)\,$, where recall $H(\rho)=-\rho\log(\rho)-(1-\rho)\log(1-\rho)$. By direct differentiation
$G_{\text{entropy}}(\theta;(x,a))=\left[\frac{1}{\sqrt{\rho(1-\rho)}}+{\lambda}\cdot\log\Big(\frac{1-\rho}{\rho}\Big)\right]\cdot\nabla_{\theta}\rho$
Following the recipe of transforming the population objective into an empirical estimate we (1) substitute population quantities with their empirical counterparts in the multiplicative factor of the gradient, (2) substitute the population gradient with the an RLOO empirical gradient. We then get
[…]
Recalling that $\widehat{G}_{\text{GRPO}}=\sqrt{\widehat{\rho}(1-\widehat{\rho})}\left$$\widehat{\nabla}_{+}-\widehat{\nabla}_{-}\right$], we find
$\widehat{G}_{\text{entropy}}(\theta;(x,a))=\left[1+{\lambda}\cdot\log\Big(\frac{1-\widehat{\rho}}{\widehat{\rho}}\Big)\cdot\sqrt{\widehat{\rho}(1-\widehat{\rho})}\right]\cdot\widehat{G}_{\text{GRPO}}\,.$
Equivalently, the advantage scores of the new method are
$A_{\text{entropy}}^{\pm}=\left[1+{\lambda}\cdot\log\Big(\frac{1-\widehat{\rho}}{\widehat{\rho}}\Big)\cdot\sqrt{\widehat{\rho}(1-\widehat{\rho})}\right]\cdot A^{\pm}\,.$
Appendix C Additional Analysis and Discussion of GRPOK
C.1 Reverse-Engineering GRPOK
Taking limit of $N\rightarrow\infty$ and replacing empirical quantities with population counterparts in Eq. (17), we find
$G_{\text{GRPO}_{K}}=\frac{1-\rho_{K}}{1-\rho}\sqrt{\rho(1-\rho)}\cdot\Delta$
where the term $\Delta=\nabla_{\theta}\rho/(\rho(1-\rho))$ is the difference of correct/wrong population gradients defined in Lemma 2. Simplifying,
$G_{\text{GRPO}_{K}}=\frac{(1-\rho)^{K-1}}{\sqrt{\rho(1-\rho)}}\cdot\nabla_{\theta}\rho\,.$
Integrating the multiplicative factor $(1-\rho)^{K-\frac{3}{2}}\rho^{-\frac{1}{2}}$ over $\rho$ the per-example surrogate reward that gives this gradient is the incomplete beta function with parameters $1/2$ and K1/2K-1/2.
${\mathcal{B}}\Big({\rho};{1/2},{K-1/2}\Big)\,.$
Written in terms of Pass@K reward, the surrogate reward is
${\mathcal{B}}\Big({1-(1-\rho_{K})^{1/K}};{1/2},{K-1/2}\Big)\,.$ (34)
By symmetry properties of the incomplete beta function, this is up to (a $K$-dependent) constant equivalent to ${\mathcal{B}}\Big({(1-\rho_{K})^{1/K}};{1/2},{K-1/2}\Big)\,.$
For $K=1$,
${\mathcal{B}}\Big({1-(1-\rho_{K})^{1/K}};{1/2},{K-1/2}\Big)={\mathcal{B}}\Big({\rho_{K}};{1/2},{1/2}\Big)=\int_{0}^{\rho_{K}}\frac{1}{\sqrt{\tau(1-\tau)}}\mathrm{d}\tau=2\arcsin(\sqrt{\rho_{K}}).$
Thus, the surrogate function coincides with that of $\widetilde{\text{GRPO}}_{K}$ (Claim 2). However, for $K\geq 2$, the two surrogates differ. Specifically, GRPOK’s surrogate is adaptive to the value of $K$. See Fig. 5.
Figure 5: A comparison of surrogate reward functions, plotted against the Pass@K reward $\rho_{K}$. The colored lines show the GRPOK surrogate, Eq. (34) for various $K$ values. The black dashed line shows the (normalized) $\widetilde{\text{GRPO}}_{K}$ surrogate, $\frac{2}{\pi}\arcsin(\sqrt{\rho_{K}})$. (All curves are normalized to reach value 11 at $\rho_{K}=1$.) The surrogates agree for $K=1$ (dashed and purple solid lines coincide), but for $K>1$ the GRPOK surrogates become concave providing stronger optimization incentive for hard examples.
C.2 GRPOK=2 vs. Skew-R
[…]
$\widehat{G}_{\text{GRPO}_{K=2}}=\frac{1-\widehat{\rho}_{2}-1/N}{1-\widehat{\rho}}\sqrt{\widehat{\rho}(1-\widehat{\rho})}\cdot\widehat{\nabla}_{+}-\frac{1-\widehat{\rho}_{2}}{1-\widehat{\rho}}\sqrt{\widehat{\rho}(1-\widehat{\rho})}\cdot\widehat{\nabla}_{-}\,$.
For $N\gg 1$, for which ${1-\widehat{\rho}_{2}-1/N}\approx 1-\widehat{\rho}_{2}\approx(1-\widehat{\rho})^{2}$, this becomes the same as the skew-R objective in Eq. (27): $(1-\widehat{\rho})\sqrt{\widehat{\rho}(1-\widehat{\rho})}\left$$\widehat{\nabla}_{+}-\widehat{\nabla}_{-}\right$]. This raises a question with regards to our reverse-engineering perspective: On the one hand, we have argued that at population level skew-R maximizes $\arcsin(\sqrt{\rho})+\sqrt{\rho(1-\rho)}$. On the other hand, we have seen in the previous subsection that GRPOK=2 maximizes ${\mathcal{B}}\Big({1-(1-\rho_{2})^{1/2}};{1/2},{3/2}\Big)$.
First, note that the two surrogates are in fact algebraically identical. To see this, we can express ${\mathcal{B}}\Big({1-(1-\rho_{2})^{1/2}};{1/2},{3/2}\Big)$ in terms of $\rho$ (using the definition $\rho_{2}=1-(1-\rho)^{2}$):
${\mathcal{B}}\Big({\rho};{1/2},{3/2}\Big)=\int_{0}^{\rho}\sqrt{\frac{1-\tau}{\tau}}\mathrm{d}\tau=\int_{0}^{\rho}\frac{1}{\sqrt{2\tau(1-\tau})}\mathrm{d}\tau+\int_{0}^{\rho}\frac{1-2\tau}{\sqrt{2\tau(1-\tau})}\mathrm{d}\tau\,.$
The first integral is $\sqrt{2}\arcsin(\sqrt{\rho})$ and the second $\sqrt{2}\sqrt{\rho(1-\rho)}$, confirming the identity. This means there is no inconsistency in our reverse-engineering strategy. At the same time, the simple three-step forward-engineering recipe operates by substituting population quantities with empirical estimates without accounting for unbiasedness (e.g., via leave-one-out): thus it cannot recover the finite-sample ($N$ not $\gg 1$) version of GRPOK and arrives at skew-R.
This algebraic identity, however, raises a subtle question of interpretation for how we talk about “regularization.” One reading (a 0/1-reward-centric lens) is to say: $\mathrm{GRPO}_{K=1}$ already optimizes the surrogate $\arcsin(\sqrt{\rho})$; moving to $K=2$ simply adds the extra term $\sqrt{\rho(1-\rho)}$, which one might then call a reward-level regularizer
However, for the Pass@K methods, we argue it is more appropriate to adopt a Pass@K-centric lens. For Pass@K methods, $\rho_{K}$ is the actual performance metric of interest, and the 0/1 reward ($\rho$) is merely an intermediate quantity used in its calculation (since $\rho_{K}=1-(1-\rho)^{K}$). From this perspective, the objective in Eq. (34) is not a ’regularized 0/1 surrogate’ but rather a direct surrogate for $\rho_{K}$: a monotone transformation of the Pass@K objective itself. Under this view, $\text{GRPO}_{K}$ is just doing standard surrogate maximization of the performance metric (analogous to cross-entropy as a surrogate for 0/1 risk), rather than “optimizing $\rho$ with an added regularizer.”
Both interpretations are mathematically consistent; the choice depends on whether one views the optimization goal as improving the 0/1 reward (and thus interpreting $K>1$ as regularization) or as directly optimizing the Pass@K metric (and thus interpreting the objective as a simple, monotone transformation of $\rho_{K}$).
C.3 Additional Related Work
In the main body, we compared $\text{GRPO}_{K}$ to its biased variant from [Mah+25] and to $\widetilde{\text{GRPO}}_{K}$ as implemented by [Che+25]. Policy optimization for Pass@K has also been studied by [WK25]. Their approach differs from both ours and [Mah+25] in two important ways.
First, [WK25] do not formulate GRPO-style estimators. The fair comparison is therefore with our $\mathrm{REINFORCE}_{K}$ / $\mathrm{RLOO}_{K}$ estimators from Sec. 3, which directly target the Pass@K objective. Like us, they construct an unbiased estimator of $\nabla_{\theta}\rho_{K,\theta}(x,a)$, but the structure of their estimator is different.
In our notation (and dropping an overall prefactor $K$ that is common across examples and matches Sec. 3), their per-example estimator for $\nabla_{\theta}\rho_{K,\theta}(x,a)$ has the following form:
$\frac{1}{N}\sum_{i\in[N]}\nabla_{\theta}\log\pi_{\theta}(y_{i}|x)\cdot\begin{cases}1,&\text{if }r_{0/1}(y_{i},a)=1,\\ 1-{\binom{N^{-}-1}{K-1}}\big/{\binom{N-1}{K-1}},&\text{if }r_{0/1}(y_{i},a)=0.\end{cases}$ (35)
This should be contrasted with our $\mathrm{REINFORCE}_{K}$ estimator which can be interpreted as a reweighting of vanilla REINFORCE. By contrast, the estimator in Eq. (35) assigns a positive coefficient not only to correct responses but also to incorrect responses. That is, it explicitly takes positive gradient steps in directions corresponding to some incorrect samples in the batch. Thus, it is also not comparable to RLOOK.
Besides, it is not immediately clear from (35) how to subtract a baseline to reduce variance while maintaining unbiasedness. Instead, going from REINFORCEK to RLOOK in Sec. 3 was rather intuitive. We refer the reader to [WK25, Sec. 4] for more discussion on variance reduction of their method.
For completeness, we can apply our reverse-engineering lens to Eq. (35) and verify that it indeed ascends the Pass@K objective in expectation. Rewriting Eq. (35) more compactly using our notation for conditional gradients $\widehat{\nabla}_{+}$ and $\widehat{\nabla}_{-}={\binom{N^{-}-1}{K-1}}\big/{\binom{N-1}{K-1}}$ and for $\widehat{f}_{K-1}^{-}$, we obtain (for a single $(x,a)$)
$\hat{\rho}\,\widehat{\nabla}_{+}\;+\;(1-\hat{\rho})\,\bigl(1-\widehat{f}_{K-1}^{-}\bigr)\,\widehat{\nabla}_{-}\,$.
Passing to population quantities (Lemmas 1 and 2) and using $\rho_{K}=1-(1-\rho)^{K}$, this expression becomes
$\rho\cdot\frac{\nabla_{\theta}\rho}{\rho}\;-\;(1-\rho)\left(1-\frac{1-\rho_{K}}{1-\rho}\right)\cdot\frac{\nabla_{\theta}\rho}{1-\rho}\;=\;\left(\frac{1-\rho_{K}}{1-\rho}\right)\nabla_{\theta}\rho\;=\;(1-\rho)^{K-1}\,\nabla_{\theta}\rho\;=\;\frac{1}{K}\,\nabla_{\theta}\rho_{K,\theta}.$
Thus, in population, the estimator from [WK25] is indeed maximizing $\rho_{K,\theta}$, consistent with our own derivations. The difference is at finite-samples: (i) how gradient weight is assigned across samples (especially incorrect ones), and (ii) how variance reduction is handled at finite $N$.
Finally, [WK25] explicitly emphasize extending Pass@K-style optimization to non-binary continuous rewards. It would be interesting future work to extend the surrogate reward lens to continuous-reward settings.